### v1

# Flyby-induced displacement: analytic solution

Abstract
Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method. Pure displacement arises when the wave zone contains an integer number of half-waves. The results confirm the prediction of Zel’dovich and Polnarev.
Key words: gravitational waves, displacement memory effect, flyby, Nikiforov-Uvarov method
pacs:
04.20.-q Classical general relativity;
I Analytic solution for flyby
Zel’dovich and Polnarev suggested that gravitational waves generated by flyby could be observed through the displacement memory effect (DM) ZelPol ; DM-1 ; Jibril . Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 . Zel’dovich - Polnarev suggestion was confirmed numerically for the derivative-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 . Their profile does not support an analytic solution, however it can be approximated by the derivative of the Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,,$ (1)
which can be studied semi-analytically using confluent Heun functions DM-1 ; DM-2 . This led us to conjecture a sort of quantisation condition: DM ariseses when the wave zone contains an integer number $m$ of half-waves. In this note we present yet another approximation of the flyby profile, namely by the Scarf potential Scarfpot ,
${\cal A}^{Scarf}(U)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (2)
for which an analytic solution is found; it confirms the above conjecture.
Note that ${\cal A}^{Scarf}(U)\propto\cosh U\,{\cal A}^{dPT}(U)$ implying that for small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly. However dilating $U$ appropriately, almost perfect overlapping is obtained, as depicted in FIG.1.
Figure 1: The (dilated) Scarf potential (2) and its Pöschl-Teller  alter-ego (1) are both good approximations of the derivative-of-Gaussian profile proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
In detail, the geodesics are, in Brinkmann coordinates, determined by the a Sturm-Liouville-type transverse equations DM-2 ,
$\dfrac{d^{2}\!X^{i}}{dU^{2}}-\frac{1}{2}{\cal A}X^{i}=0\,,$ (3)
$i=1,2$, where $U$ behaves as non-relativistic time and ${\cal A}={\cal A}(U)$ is the profile of the gravitational wave DBKP ; DGH91 . These equations are completed with the DM boundary conditions,
$\frac{d{\bm{X}}}{dU}\Big{|}(U=-\infty)=0=\frac{d{\bm{X}}}{dU}\Big{|}(U=+\infty)\,.$ (4)
Below we show that the hyperbolic Scarf potential Scarfpot ; Scrf1 , eqn. (2), allows us to find an analytic solution which confirms the quantisation conjecture above. Eqn.(3) can be viewed indeed as a Schrödinger equation for the (non-normalizable) zero-energy ground state DM-1 . The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile $(g/2)\cosh^{-2}U$ into $(1-t^{2})\frac{d^{2}X}{dt^{2}}-2t\frac{dX}{dt}-\frac{g}{4}X=0\,,$ which can be solved by Legendre polynomials DM-1 .
$(1+t^{2})\frac{d^{2}X}{dt^{2}}+t\frac{dX}{dt}-\frac{gt}{1+t^{2}}X=0\,,$ (5)
which is of the generalised hypergeometric type, and can be solved analytically by the Nikiforov-Uvarov method NikiUvar . Since a fairly detailed description of this simple, yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , we only provide a brief summary of the algorithm in the Appendix.
The Nikoforov-Uvarov method (recalled in the Appendix) applies to second-order differential equations of the generalised hypergeometric type of the form,
$u^{\prime\prime}+\frac{\pi_{1}(z)}{\sigma(z)}u^{\prime}+\frac{\sigma_{1}(z)}{\sigma^{2}(z)}u=0\,.$ (6)
For Eq. (5) in particular,
$\sigma=1+t^{2},\;\;\;\pi_{1}=t,\;\;\;\sigma_{1}=-gt\,.$ (7)
Then,
$\sigma_{3}=\frac{t^{2}}{4}+gt+k(1+t^{2})=t^{2}\left(k+\frac{1}{4}\right)+gt+k\,,$ (8)
Therefore, the condition $\det(\sigma_{3})=g^{2}-4k\left(k+\frac{1}{4}\right)=0$ gives
$k=\frac{\pm\sqrt{1+16g^{2}}-1}{8},$ (9)
and
$|g|=2\sqrt{k\left(k+\frac{1}{4}\right)}\,.$ (10)
Thus we have
$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.$ (11)
Then
$\tau=\pi_{1}+2\pi=2t\pm\frac{|g|}{\sqrt{k}}\left|t+\frac{2k}{g}\right|.$ (12)
To have real solutions, we assume $k>0$, that is we choose upper sign in Eq.(9): $k=\frac{\sqrt{1+16g^{2}}-1}{8}.$ On the other side, the quantization condition Eq.(35) requires
$\lambda=-n(n+1)\mp\epsilon\,\frac{n|g|}{\sqrt{k}},$ (13)
since $\tau^{\prime}=2\pm\epsilon\,\frac{|g|}{\sqrt{k}}$ with $\epsilon=\mathrm{sign}\left(t+\frac{2k}{g}\right)\,.$ Therefore we get, after some algebra,
$k=\lambda-\pi^{\prime}=-n^{2}-n-\frac{1}{2}\mp\epsilon\frac{|g|}{\sqrt{k}}\left(n+\frac{1}{2}\right),$
and using Eq. (10) for $|g|/\sqrt{k}$, we get
$\sqrt{k+\frac{1}{4}}=\mp\epsilon\left(n+\frac{1}{2}\right)$ (14)
and we see that we must choose the lower signs in Eq.(11) and Eq.(12) to get $\mp\epsilon=1$. Then the solution of Eq.(14) is
$k_{n}=n(n+1),$ (15)
and Eq.(10) determines the “magic” strengths of the Scarf potential Eq.(2), for which we have polynomial solutions for $y_{n}(t)$ labeled by the integer “quantum” number $n$,
$\;\;\displaystyle{|g_{n}|=(2n+1)\sqrt{n(n+1)}\;.}\;\;$ (16)
It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. Because of Eq.(15) and Eq.(16), we have,
$\pi(t)=-nt-\mathrm{sign}(g)\sqrt{n(n+1)},\;\;\;\tau(t)=-(2n-1)t-2\mathrm{sign}(g)\sqrt{n(n+1)}\,.$ (17)
The equation for the gauge function,
$\frac{d\varphi}{dt}=-\frac{nt+\mathrm{sign}(g)\sqrt{n(n+1)}}{1+t^{2}}\,,$ (18)
can be easily integrated and we get (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
$\varphi(t)=-\frac{n}{2}\ln(1+t^{2})+\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}\,.$ (19)
Therefore, the corresponding geodesic trajectory is given by
$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,.}\;\;$ (20)
The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq.(36) in which we need to find the weight function $\rho(t)$, for which we have the equation,
$\frac{((1+t^{2})\rho)^{\prime}}{(1+t^{2})\rho}=\frac{\tau}{\sigma}=-\frac{(2n-1)t+2\,\mathrm{sign}(g)\sqrt{n(n+1)}}{1+t^{2}}\,,$ (21)
which can be again easily integrated with the result,
$\rho(t)=(1+t^{2})^{-\left(n+\frac{1}{2}\right)}e^{-2\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,.$ (22)
Therefore
$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,.}\;\;$ (23)
The normalization constants $B_{n}$ should be determined from the initial conditions. The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g)$, $y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23$ and $y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)$. These polynomials can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as follows: $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$, with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
For (half)wave numbers $n=1,2,3,4$ the analytic solutions (20) reproduce almost identically those half-numerical ones shown in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case. For $n\geq 5$, though, DM requires to find a high number of decimals and the numerical procedure becomes tedious. Plotting instead the analytic solution (20)-(23) yields FIG.2.
Figure 2: DM geodesics for the Scarf profile (2) for wave numbers $n=5$ and $n=6$.
In conclusion, our results show that the profile proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by the scarf potential (2) in FIG.1 which supports analytic solutions, (20)-(23), labeled by an integer $n$, (16), which can be constructed by the Nikiforov-Uvarov method. The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.2, extending an observation made DM-2 for the Gaussian/Pöschl-Teller profile.
Further physical applications of the Scarf potential which include supersymmetry Scarfpot ; Compean , non-central potentials Dutt and even quarks (QCD), are reviewed in ref. Raposo . Note that, in the supersymmetry literature, the Scarf hyperbolic potential, which has recently attracted much attention Scrf1 , was originally called the Gendenstein potential Natanson .
Acknowledgements.
The authors are grateful to Gary Gibbons for discussions. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320).
Appendix A Compendium on the Nikoforov-Uvarov method
As it was already mentioned in the text, the Nikoforov-Uvarov method applies to second-order differential equations of the generalised hypergeometric type of the form
$u^{\prime\prime}+\frac{\pi_{1}(z)}{\sigma(z)}u^{\prime}+\frac{\sigma_{1}(z)}{\sigma^{2}(z)}u=0\,,$ (24)
where the prime denotes $d/dz$. $z$ can be complex. $\pi_{1}(z)$ is a polynomial at most of the first degree, and $\sigma(z)$, $\sigma_{1}(z)$ are polynomials at most of the second degree NikiUvar .
The set of solutions of Eq. (24) is invariant under “gauge” transformations $u(z)\to y(z)$, $u(z)=e^{\varphi(z)}y(z),$ if the gauge function $\varphi(z)$ satisfies the equation
$\varphi^{\prime}(z)=\frac{\pi(z)}{\sigma(z)}\,$ (25)
where $\pi(z)$ is a polynomial at most of the first degree. In this case for $y(z)$ we obtain also a generalised hypergeometric-type equation:
$y^{\prime\prime}+\frac{\pi_{2}(z)}{\sigma(z)}y^{\prime}+\frac{\sigma_{2}(z)}{\sigma^{2}(z)}y=0\,,$ (26)
where $\pi_{2}(z)=\pi_{1}(z)+2\pi(z)$ is a polynomial of at most of the first degree, and
$\sigma_{2}(z)=\sigma_{1}(z)+\pi^{2}(z)+\pi(z)\left[\pi_{1}(z)-\sigma^{\prime}(z)\right]+\pi^{\prime}(z)\sigma(z)$ (27)
is a polynomial of at most of the second degree. Choosing $\pi(z)$ so that
$\sigma_{2}(z)=\lambda\sigma(z)\,$ (28)
where $\lambda$ is a constant, Eq. (6) is simplified to an hypergeometric type equation,
$\sigma(z)y^{\prime\prime}+\pi_{2}(z)y^{\prime}+\lambda y=0\,.$ (29)
In the light of Eq. (27), Eq. (28) means
$\pi^{2}+\pi[\pi_{1}-\sigma^{\prime}]+\sigma_{1}-k\sigma=0,$ (30)
where $k=\lambda-\pi^{\prime}$ is another constant. Eq. (30) is solved by,
$\pi=\frac{\sigma^{\prime}-\pi_{1}}{2}\pm\sqrt{\sigma_{3}(z)}{\quad\text{\small where}\quad}\sigma_{3}(z)=\left(\frac{\sigma^{\prime}-\pi_{1}}{2}\right)^{2}-\sigma_{1}+k\sigma\,.$ (31)
$\pi$ is a polynomial only when $\sigma_{3}(z)$ is the square of a first order polynomial. Hence it has a double root and its discriminant is zero, $\Delta(\sigma_{3})=0\,,$ which determines the constant $k$, and hence the constant $\lambda$.
We want to find polynomial solutions of Eq. (29). It can be shown that derivatives of $y(z)$, $v_{n}(z)=y^{(n)}(z)$, are also generalised hypergeometric-type functions,
$\sigma v_{n}^{\prime\prime}+\tau_{n}(z)v_{n}^{\prime}+\mu_{n}v_{n}=0\,,$ (32)
and we get the recurrence relations
$\tau_{n}(z)=\sigma^{\prime}(z)+\tau_{n-1}(z),\;\;\;\;\mu_{n}=\mu_{n-1}+\tau_{n-1}^{\prime}\,,$ (33)
with initial values $\tau(z)=\tau_{0}(z)=\pi_{2}(z),\;\mu_{0}=\lambda\,.$ If $y(z)=y_{n}(z)$ is a polynomial of order $n$, then $v_{n}=\mathrm{const}$ and Eq. (32) is satisfied only if $\mu_{n}=0$. But the repeated application of the recurrence relations Eq. (33) produces
$\tau_{n}(z)=n\sigma^{\prime}(z)+\tau(z),\;\;\;\mu_{n}=\lambda+n\tau^{\prime}+\frac{1}{2}n(n-1)\sigma^{\prime\prime}\,.$ (34)
Getting a polynomial solution $y(z)=y_{n}(z)$ of Eq. (29) thus requires the “quantization condition”,
$\lambda=\lambda_{n}=-n\tau^{\prime}-\frac{1}{2}n(n-1)\sigma^{\prime\prime}.$ (35)
It can be shown NikiUvar that, when the quantization condition Eq. (35) is satisfied, the corresponding polynomial solution of Eq. (29) is given by the Rodrigues formula
$y_{n}(z)=\frac{B_{n}}{\rho(z)}\left[\sigma^{n}(z)\rho(z)\right]^{(n)},$ (36)
where $B_{n}$ is some (normalization) constant, and the weight functions $\rho(z)$ satisfies Pearson’s equation Ismail
$(\sigma\rho)^{\prime}=\rho\tau\,.$ (37)

### v2

@@
 Abstract
-Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method.
+Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the hyperbolic Scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method applied to the Scarf potential.
 Pure displacement arises when the wave zone contains an integer number of half-waves.
@@
 04.20.-q Classical general relativity;
+preprint: arXiv: 2502.0136v2 [gr-qc]
 I Analytic solution for flyby
@@
 Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 .
-Zel’dovich - Polnarev suggestion was confirmed numerically for the derivative-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 .
-Their profile does not support an analytic solution, however it can be approximated by the derivative of the Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
+Zel’dovich - Polnarev suggestion was confirmed numerically for the derived-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 .
+Their profile does not support an analytic solution, however it can be approximated by the derived Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
 ${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,,$ (1)
@@
 This led us to conjecture a sort of quantisation condition: DM ariseses when the wave zone contains an integer number $m$ of half-waves.
-In this note we present yet another approximation of the flyby profile, namely by the Scarf potential Scarfpot ,
+In this note we present yet another approximation of the flyby profile, namely by the hyperbolic Scarf potential scarfpot ,
 ${\cal A}^{scarf}(U)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (2)
 for which an analytic solution is found; it confirms the above conjecture.
-Note that ${\cal A}^{Scarf}(U)\propto\cosh U\,{\cal A}^{dPT}(U)$ implying that for small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly.
-However dilating $U$ appropriately, almost perfect overlapping is obtained, as depicted in FIG.1.
-Figure 1: The (dilated) Scarf potential (2) and its Pöschl-Teller  alter-ego (1) are both good approximations of the derivative-of-Gaussian profile proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
+For small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1.
+Figure 1: The derived Pöschl-Teller  profile (1) and its Scarf potential alter-ego (2) are both good approximations of the derived-Gaussian one proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
 In detail, the geodesics are, in Brinkmann coordinates, determined by the a Sturm-Liouville-type transverse equations DM-2 ,
@@
 Eqn.(3) can be viewed indeed as a Schrödinger equation for the (non-normalizable) zero-energy ground state DM-1 .
-The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile $(g/2)\cosh^{-2}U$ into $(1-t^{2})\frac{d^{2}X}{dt^{2}}-2t\frac{dX}{dt}-\frac{g}{4}X=0\,,$ which can be solved by Legendre polynomials DM-1 .
+The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ into a form which can be solved by Legendre polynomials DM-1 .
 $(1+t^{2})\frac{d^{2}X}{dt^{2}}+t\frac{dX}{dt}-\frac{gt}{1+t^{2}}X=0\,,$ (5)
@@
 $\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,.}\;\;$ (20)
-The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq.(36) in which we need to find the weight function $\rho(t)$, for which we have the equation,
-$\frac{((1+t^{2})\rho)^{\prime}}{(1+t^{2})\rho}=\frac{\tau}{\sigma}=-\frac{(2n-1)t+2\,\mathrm{sign}(g)\sqrt{n(n+1)}}{1+t^{2}}\,,$ (21)
-which can be again easily integrated with the result,
+The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (35) in which we need to find the weight function $\rho(t)$, for which we found,
 $\rho(t)=(1+t^{2})^{-\left(n+\frac{1}{2}\right)}e^{-2\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,.$ (21)
@@
 The normalization constants $B_{n}$ should be determined from the initial conditions.
-The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g)$, $y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23$ and $y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)$.
-These polynomials can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as follows: $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$, with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
+The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$ They can be expressed in terms of Routh-Romanovski polynomials scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
 For (half)wave numbers $n=1,2,3,4$ the analytic solutions (20) reproduce almost identically those half-numerical ones shown in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case.
@@
 Raposo .
-Note that, in the supersymmetry literature, the Scarf hyperbolic potential, which has recently attracted much attention Scrf1 , was originally called the Gendenstein potential Natanson .
+Note that, in the supersymmetry literature, the Scarf hyperbolic potential has recently attracted much attention Scrf1 .
+This potential was discovered by Gendenshtein and is therefore referred also to as the Gendenstein potential Gendenon ; Natanson .
+See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for related work.
+For isotropic profile we have a “universal model” : every plane gravitational wave can be mapped locally conformally to free Minkowski space by following Niederer, or Arnold Niederer ; Arnold ; Takagi , consistently with the vanishing of the Cotton or Weyl tensor Inomata ; Dhasmana ; GibbonsDark ; ZZH21 .
+However gravitational waves are isotropic only in 1d, and their anisotropy in d=2 dim implies that we arrive to unchartered territory Lukash ; Inomata ; ZZH21 and it is not clear whether we still have a (conformal) “universal ancestor”.
 Acknowledgements.

### v3

@@
-Flyby-induced displacement: analytic solution
+Flyby-induced displacement effect: an analytic solution
 Abstract
-Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the hyperbolic Scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method applied to the Scarf potential.
-Pure displacement arises when the wave zone contains an integer number of half-waves.
-The results confirm the prediction of Zel’dovich and Polnarev.
-Key words: gravitational waves, displacement memory effect, flyby, Nikiforov-Uvarov method
+We describe analytically the scattering of particles by a sandwich gravitational wave generated during a flyby.
+The derivative-of-the-Gaussian profile proposed by Gibbons and Hawking is approximated by the hyperbolic scarf potential, which allows for an exact analytic solution via the Nikiforov-Uvarov method.
+Our results confirm the prediction of Zel’dovich and Polnarev about certain “magical” amplitudes of the potential.
+Key words: gravitational waves; displacement memory effect; flyby; Nikiforov-Uvarov method.
 pacs:
 04.20.-q Classical general relativity;
-preprint: arXiv: 2502.0136v2 [gr-qc]
-I Analytic solution for flyby
-Zel’dovich and Polnarev suggested that gravitational waves generated by flyby could be observed through the displacement memory effect (DM) ZelPol ; DM-1 ; Jibril .
-Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 .
-Zel’dovich - Polnarev suggestion was confirmed numerically for the derived-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 .
-Their profile does not support an analytic solution, however it can be approximated by the derived Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
-${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,,$ (1)
-which can be studied semi-analytically using confluent Heun functions DM-1 ; DM-2 .
-This led us to conjecture a sort of quantisation condition: DM ariseses when the wave zone contains an integer number $m$ of half-waves.
-In this note we present yet another approximation of the flyby profile, namely by the hyperbolic Scarf potential scarfpot ,
-${\cal A}^{scarf}(U)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (2)
-for which an analytic solution is found; it confirms the above conjecture.
-For small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1.
+preprint: arXiv: 2502.0136v3 [gr-qc]
+I Introduction
+One of the early proposals to observe gravitational waves , put forward by Braginsky and Thorne BraTho , was to study the motion of particles initially in rest after the passing of a burst of gravitational waves .
+They argued indeed that Burst With Memory could be among the earliest kinds of gravitational waves detected BraTho .
+Initial studies Ehlers considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories.
+Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : …although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.
+The GW memory effect offers a way to test general relativity including the nonlinear nature of gravity, and provides additional information about the GW source.
+Preliminary studies have suggested that the detection of GW memory from individual mergers of massive binary black holes is expected with LISA Favata .
+However, more realistic estimates show that observing the GW memory effect at detectors such as LISA, although feasible, is a challenging task fly-1 ; fly-2 .
+Recent attempts to detect the effect in LIGO/Virgo’s first transient gravitational-wave catalog, as well as in a pulsar timing array experiment NANOGrav, have predictably yielded negative results fly-3 ; fly-4 .
+However the prospects for the next generation of gravitational wave detectors are brighter even for future ground-based detectors such as Cosmic Explorer and the Einstein Telescope (see fly-2 ; fly-4 and references therein).
+Zel’dovich and Polnarev did not present any proof.
+Although their suggestion could be confirmed in Christo ; BMS1 ; BMS2 ; BMS3 ; BMS4 , there were some doubts due to the apparent contradiction with the Bondi-Pirani theorem on the caustic property of plane gravitational waves BoPi89 ; ShortMemory .
+Then it came as a surprise that that the DM effect could be confirmed numerically DM-1 ; Jibril ; DM-2 for plane waves — provided the amplitude takes some “magical” values.
+Below we present an analytic proof.
+In Brinkmann coordinates, the metric is written as Brinkmann ,
+$ds^{2}=d{\bm{X}}^{2}+2dUdV+{\cal A}(U)\big{(}X_{+}^{2}-X_{-}^{2}\big{)}dU^{2}\,.$ (1)
+Gibbons and Hawking GibbHaw71 argued that in linear theory the Riemann tensor of a gravitational wave is proportional to the fourth time derivative of the quadrupole moment of the source GibbonsP .
+This lead them to propose the derivative of the Gaussian,
+${\cal A}^{dG}\propto\frac{\;d}{dU}\Big{(}\exp[-U^{2}]\Big{)}\,.$ (2)
+as profile.
+The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
+${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,.$ (3)
+Then the geodesics can be studied semi-analytically DM-1 ; DM-2 .
+Our investigations led us to conjecture that DM arises when a quantization holds, namely when the wave zone contains an integer number of what we called half-waves.
+In this note we present a third approximation of the flyby profile (2), namely by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was first introduced in Scarf to mimic the potential of a real crystal.
+It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt1 .,
+${\cal A}^{Scarf}(U)=2g\frac{\;\;d}{dU}\left(\frac{1}{\cosh U}\right)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (4)
+to be compared with (3).
+For small values of $U$ the Scarf profile (4) behaves as the derivative of the Gaussian or of Pöschl-Teller , (2) or (3), but for large $U$ it falls off more slowly.
+However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1.
+Their similarity follows from that all three of them are derivatives of similar bell-shaped functions.
+We note also all three profiles in (2)-(3)-(4) are derivatives of an even function and are therefore odd.
 Figure 1: The derived Pöschl-Teller  profile (3) and its Scarf potential alter-ego (4) are both good approximations of the derived-Gaussian one proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
-In detail, the geodesics are, in Brinkmann coordinates, determined by the a Sturm-Liouville-type transverse equations DM-2 ,
-$\dfrac{d^{2}\!X^{i}}{dU^{2}}-\frac{1}{2}{\cal A}X^{i}=0\,,$ (3)
-$i=1,2$, where $U$ behaves as non-relativistic time and ${\cal A}={\cal A}(U)$ is the profile of the gravitational wave DBKP ; DGH91 . These equations are completed with the DM boundary conditions,
+The geodesics are, in Brinkmann coordinates, determined by the transverse equations DM-2 ,
+$\dfrac{d^{2}\!X^{\pm}}{dU^{2}}\mp\frac{1}{2}{\cal A}X^{\pm}=0\,,$ (5)
+where the $\pm$ label refers to the upper resp.
+lower component.
+The sign of the second  term comes from the relative minus between the coordinates in (1).
+The equation for the $V$ component follows from these DM-1 ; DM-2 , and is therefore omitted.
+We recall that $U$ behaves, in the Bargmann framework DBKP ; DGH91 , as non-relativistic time.
+These Sturm-Liouville-type equations are completed with the DM boundary conditions,
 $\frac{d{\bm{X}}}{dU}\Big{|}(U=-\infty)=0=\frac{d{\bm{X}}}{dU}\Big{|}(U=+\infty)\,.$ (6)
-Below we show that the hyperbolic Scarf potential scarfpot ; Scrf1 , eqn.
-(2), allows us to find an analytic solution which confirms the quantisation conjecture above.
-Eqn.(3) can be viewed indeed as a Schrödinger equation for the (non-normalizable) zero-energy ground state DM-1 .
-The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ into a form which can be solved by Legendre polynomials DM-1 .
-$(1+t^{2})\frac{d^{2}X}{dt^{2}}+t\frac{dX}{dt}-\frac{gt}{1+t^{2}}X=0\,,$ (5)
-which is of the generalised hypergeometric type, and can be solved analytically by the Nikiforov-Uvarov method NikiUvar .
-Since a fairly detailed description of this simple, yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , we only provide a brief summary of the algorithm in the Appendix.
+Our main result then says that for the profile (4) analytic solutions can be found which confirm the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above.
+Our clue is that Eqn.(5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its (non-normalizable) zero-energy ground state solutions DM-1 .
+II Analytic solution for flyby
+We first recall how things go for the simple Pöschl-Teller case. DM-1 ; DM-2 . Putting $t=\tanh U$, into eqn. (5) with profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ yields
+$(1-t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}-2t\frac{dX^{\pm}}{dt}\mp\frac{g}{4}X^{\pm}=0\,,$ (7)
+Let us consider first one of the components, say $X^{+}$. Then DM is obtained in terms of Legendre polynomials DM-1 when the coefficient in the linear-in-$X^{+}$ term is
+$g=-4n(n+1){\quad\text{\small where}\quad}n>0\quad\text{is a natural integer}\,.$ (8)
+However passing to the other component, the linear-in-$X^{-}$ term has opposite sign, therefore DM would be obtained for $-g$.
+But this is impossible because $g$ has already been fixed.
+In conclusion, one has what we called “half DM” DM-2 : we get a DM solution in one or in the other sector, but not for both in general — unless the “wrong” component is set to zero.
+Solutions for the Scarf potential ${\cal A}^{Scarf}$ (and similarly for derived Pöschl-Teller , ${\cal A}^{dPT}$), DM-2 ), are obtained by putting $t=\sinh{U}$. Then (5) is written
+$(1+t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}+t\frac{dX^{\pm}}{dt}{\mp}g\frac{t}{1+t^{2}}X^{\pm}=0\,.$ (9)
+The two equations here are again uncoupled and can therefore be solved independently.
+Note however the coefficient $t$ in front of the linear-in-$X^{\pm}$ term which compensates the $\pm$ sign change between the components.
+Therefore the same $g$ works for both components, as it will be seen in sec.
+III.
+Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller .
+For derived Pöschl-Teller it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions in eqn.
+# (IV.16) of DM-2 .
+Turning to Scarf, we note that (9) is of the generalised hypergeometric type which can be solved analytically by the Nikiforov-Uvarov (NU) method NikiUvar .
+Since a fairly detailed description of this simple yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , here we just list the main result.
+The algorithm, whose details are discussed in sec.III, yields the desired “magic” coefficient $g$ in (9),
+$|g|=|g_{n}|=(2n+1)\sqrt{n(n+1)}\,,$ (10)
+where $n$ is an integer. Then the trajectory is,
+$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,,}\;\;$ (11)
+where $y_{n}(t)$ is the polynomial
+$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,}\;\;$ (12)
+with weight function
+$\rho(t)=(1+t^{2})^{-\left(n+\frac{1}{2}\right)}e^{-2\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,.$ (13)
+The normalization constants $B_{n}$ is determined from the initial conditions.
+The analytic solutions (11) reproduce almost identically those obtained numerically in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case.
+For $n\geq 5$, though, the numerical procedure becomes tedious as it requires to calculate a high number of decimals.
+For high wave numbers the result becomes also inaccurate because the procedure used in DM-2 is only an approximation.
+Plotting instead the analytic solution (11)-(12), FIG.3, is straightforward, underlining the advantage of the analytical solution.
+Figure 2: The DM geodesics for the derived Pöschl-Teller (3) and for the Scarf profiles (4) are similar however not fully identical, as shown for wave numbers ${\bf n=1}$ and ${\bf n=2}$.
+(The trajectories were rescaled for better comparison).
+Figure 3: DM geodesics for the Scarf profile (4) with wave numbers ${\bf n=5}$ and ${\bf n=6}$.
+III Transverse geodesics by the Nikiforov-Uvarov method
 The Nikoforov-Uvarov method (recalled in the Appendix) applies to second-order differential equations of the generalised hypergeometric type of the form,
 $u^{\prime\prime}+\frac{\pi_{1}(z)}{\sigma(z)}u^{\prime}+\frac{\sigma_{1}(z)}{\sigma^{2}(z)}u=0\,.$ (14)
-For Eq. (5) in particular,
-$\sigma=1+t^{2},\;\;\;\pi_{1}=t,\;\;\;\sigma_{1}=-gt\,.$ (7)
-Then,
-$\sigma_{3}=\frac{t^{2}}{4}+gt+k(1+t^{2})=t^{2}\left(k+\frac{1}{4}\right)+gt+k\,,$ (8)
+In the Scarp case Eq. (9) $\sigma=1+t^{2},\;\pi_{1}=t,\;\sigma_{1}=-gt\,.$ Then,
+$\sigma_{3}=\frac{t^{2}}{4}+gt+k(1+t^{2})=t^{2}\left(k+\frac{1}{4}\right)+gt+k\,.$ (15)
 Therefore, the condition $\det(\sigma_{3})=g^{2}-4k\left(k+\frac{1}{4}\right)=0$ gives
-$k=\frac{\pm\sqrt{1+16g^{2}}-1}{8},$ (9)
-and
-$|g|=2\sqrt{k\left(k+\frac{1}{4}\right)}\,.$ (10)
+$k=\frac{\pm\sqrt{1+16g^{2}}-1}{8}{\quad\text{\small and}\quad}|g|=2\sqrt{k\left(k+\frac{1}{4}\right)}\,.$ (16)
 Thus we have
-$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.$ (11)
+$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.\quad$ (17)
 Then
 $\tau=\pi_{1}+2\pi=2t\pm\frac{|g|}{\sqrt{k}}\left|t+\frac{2k}{g}\right|.$ (18)
-To have real solutions, we assume $k>0$, that is we choose upper sign in Eq.(9): $k=\frac{\sqrt{1+16g^{2}}-1}{8}.$ On the other side, the quantization condition Eq.(34) requires
+To have real solutions, we assume $k>0$, that is we choose th upper sign $k=\frac{\sqrt{1+16g^{2}}-1}{8}$ in Eq.(16). On the other side, the quantization condition Eq. (36) requires
 $\lambda=-n(n+1)\mp\epsilon\,\frac{n|g|}{\sqrt{k}}\,,$ (19)
-since $\tau^{\prime}=2\pm\epsilon\,\frac{|g|}{\sqrt{k}}$ with $\epsilon=\mathrm{sign}\left(t+\frac{2k}{g}\right)\,.$ Therefore we get, after some algebra,
+since $\tau^{\prime}=2\pm\epsilon\,\frac{|g|}{\sqrt{k}}$ with $\epsilon=\mathrm{sign}\left(t+\frac{2k}{g}\right)\,$ and we get, after some algebra,
 $k=\lambda-\pi^{\prime}=-n^{2}-n-\frac{1}{2}\mp\epsilon\frac{|g|}{\sqrt{k}}\left(n+\frac{1}{2}\right),$
-and using Eq. (10) for $|g|/\sqrt{k}$, we get
-$\sqrt{k+\frac{1}{4}}=\mp\epsilon\left(n+\frac{1}{2}\right)$ (14)
-and we see that we must choose the lower signs in Eq.(11) and Eq.(12) to get $\mp\epsilon=1$. Then the solution of Eq.(14) is
-$k_{n}=n(n+1),$ (15)
-and Eq.(10) determines the “magic” strengths of the Scarf potential Eq.(2), for which we have polynomial solutions for $y_{n}(t)$ labeled by the integer “quantum” number $n$,
-$\;\;\displaystyle{|g_{n}|=(2n+1)\sqrt{n(n+1)}\;.}\;\;$ (16)
-It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. Because of Eq. (15) and Eq. (16), we have,
+and using Eq.
+(16) for $|g|/\sqrt{k}$, we get $\sqrt{k+\frac{1}{4}}=\mp\epsilon\left(n+\frac{1}{2}\right)$ and we see that we must choose the lower signs in Eq.(17) and Eq.(18) to get $\mp\epsilon=1$.
+Then the solution is $k_{n}=n(n+1)\,,$ and Eq.(16) determines the “magic” strengths labeled by the integer “quantum” number $n$, eqn.
+(10), for which we have polynomial solutions $y_{n}(t)$.
+The $y_{n}(t)$ can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
+The first three of them are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$
+It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. We have,
 $\pi(t)=-nt-\mathrm{sign}(g)\sqrt{n(n+1)},\;\;\;\;\tau(t)=-(2n-1)t-2\mathrm{sign}(g)\sqrt{n(n+1)}\,.$ (20)
@@
 $\frac{d\varphi}{dt}=-\frac{nt+\mathrm{sign}(g)\sqrt{n(n+1)}}{1+t^{2}}\,,$ (21)
-can be easily integrated and we get (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
+can be easily integrated to yield (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
 $\varphi(t)=-\frac{n}{2}\ln(1+t^{2})+\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}\,.$ (22)
-Therefore, the corresponding geodesic trajectory is given by
-$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,.}\;\;$ (20)
-The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (35) in which we need to find the weight function $\rho(t)$, for which we found,
-$\rho(t)=(1+t^{2})^{-\left(n+\frac{1}{2}\right)}e^{-2\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,.$ (21)
-Therefore
-$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,.}\;\;$ (22)
-The normalization constants $B_{n}$ should be determined from the initial conditions.
-The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$ They can be expressed in terms of Routh-Romanovski polynomials scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
-For (half)wave numbers $n=1,2,3,4$ the analytic solutions (20) reproduce almost identically those half-numerical ones shown in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case.
-For $n\geq 5$, though, DM requires to find a high number of decimals and the numerical procedure becomes tedious.
-Plotting instead the analytic solution (20)-(22) yields FIG.2.
-Figure 2: DM geodesics for the Scarf profile (2) for wave numbers $n=5$ and $n=6$.
-In conclusion, our results show that the profile proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by the Scarf potential (2) in FIG.1 which supports analytic solutions, (20)-(22), labeled by an integer $n$, (16), which can be constructed by the Nikiforov-Uvarov method.
-The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.2, extending an observation made DM-2 for the Gaussian/Pöschl-Teller profile.
-Further physical applications of the Scarf potential which include supersymmetry scarfpot ; Compean , non-central potentials Dutt and even quarks (QCD), are reviewed in ref.
+Therefore, the corresponding geodesic trajectory is given by eqn.
+(11).
+The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq.
+(37) which requires to calculate the weight function $\rho(t)$ in (13).
+Then we end up with eqn.
+(12).
+Eqn.
+(13) implies that $\rho(t)$ does not change when $t$ and $g$ change signs simultaneously, while due to the derivatives in eqn.
+(12), the polynomials $y_{n}(t)$ acquire a $(-1)^{n}$ factor.
+Then eqn.
+(11) implies the parity behaviour
+$X^{\pm}(-t)=(-1)^{n}X^{\mp}(t)\,$ (23)
+mentioned in sec.II. This curious pairing of the components is reminiscent of and is indeed closely related to supersymmetry MSUSY .
+IV Conclusion
+The profile (2) proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by both the derived-Pöschl-Teller (3) and by the Scarf potentials (4), as illustrated in FIG.1.
+The advantage of Scarf is that it supports analytic solutions, (11)-(12), labeled by an integer $n$, eqn (10).
+They are constructed by the Nikiforov-Uvarov method.
+The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.3, extending the observation made in DM-2 for the Gaussian and Pöschl-Teller profiles.
+The origin of half-waves is clear: their number is that of the zeros of the polynomial $y_{n}(t)$.
+Our Scarf solution is similar to but slightly different from those found semi-analytically for derived-Pöschl-Teller DM-2 , as seen in FIG.2.
+It has also the advantage to be exact, while the one in DM-2 is only approximate, with error increasing for higher wave numbers.
+The behavior of both the Scarf (4) and of the derived-Pöschl-Teller (3) is explained by their parity behaviour, (23), which explains in turn why do we get DM for both components.
+The same is true for any odd-order derivative of an even-profile function DM-1 ; DM-2 ; Benin .
+For even-order derivatives the analogous choices of the parameters would be contradictory, implying that one can gets DM only for one, but not for both components — unless the other is put to zero.
+Physical applications of various incarnations of the Scarf potential are reviewed in ref.
 Raposo .
-Note that, in the supersymmetry literature, the Scarf hyperbolic potential has recently attracted much attention Scrf1 .
-This potential was discovered by Gendenshtein and is therefore referred also to as the Gendenstein potential Gendenon ; Natanson .
-See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for related work.
-For isotropic profile we have a “universal model” : every plane gravitational wave can be mapped locally conformally to free Minkowski space by following Niederer, or Arnold Niederer ; Arnold ; Takagi , consistently with the vanishing of the Cotton or Weyl tensor Inomata ; Dhasmana ; GibbonsDark ; ZZH21 .
-However gravitational waves are isotropic only in 1d, and their anisotropy in d=2 dim implies that we arrive to unchartered territory Lukash ; Inomata ; ZZH21 and it is not clear whether we still have a (conformal) “universal ancestor”.
+They include supersymmetry Scarfpot ; Scrf1 ; Compean ; Gendenshtein ; MSUSY and non-central potentials Dutt .
+The recent paper MSUSY , which is directly relevant for our purposes here, highlights the prominent role of SUSY for gravitational waves with Scarf profile.
+See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for further related work.
+We mention that our main tool here, namely the Nikiforov-Uvarov method, applies also to a number of other singular and periodic GW profiles, such as
+${\cal A}=2g\left(1-\frac{1}{\tanh^{2}{U}}\right),\;\>{\cal A}=-2g\operatorname{cn}^{2}{\left(U,\frac{1}{\sqrt{2}}\right)},\;\;{\cal A}=2g\tan{U},\;\;{\cal A}=\frac{2g}{\tan{U}}\,,$ (24)
+where $\operatorname{cn}(z,k)$ is Jacobian elliptic cosine function.
+Interestingly, the last profile corresponds to the Coulomb potential in constant curvature space $\mathbb{S}^{3}$, and hydrogenlike atoms in constant curvature spaces are most naturally considered by the Nikiforov-Uvarov method Alizzi .
 Acknowledgements.
-The authors are grateful to Gary Gibbons for discussions. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320).
+The authors are grateful to Gary Gibbons and to M.
+Elbistan for discussions and to G.
+Natanson for correspondance.
+PMZ was partially supported by the National Natural Science Foundation of China (Grant No.
+11975320).
 Appendix A Compendium on the Nikoforov-Uvarov method
@@
 $u^{\prime\prime}+\frac{\pi_{1}(z)}{\sigma(z)}u^{\prime}+\frac{\sigma_{1}(z)}{\sigma^{2}(z)}u=0\,,$ (25)
-where the prime denotes $d/dz$. $z$ can be complex. $\pi_{1}(z)$ is a polynomial at most of the first degree, and $\sigma(z)$, $\sigma_{1}(z)$ are polynomials at most of the second degree NikiUvar .
+where the prime denotes $d/dz$ where $z$ can be complex.
+$\pi_{1}(z)$ is a polynomial at most of the first degree, and $\sigma(z)$, $\sigma_{1}(z)$ are polynomials at most of the second degree NikiUvar .
 The set of solutions of Eq. (25) is invariant under “gauge” transformations $u(z)\to y(z)$, $u(z)=e^{\varphi(z)}y(z),$ if the gauge function $\varphi(z)$ satisfies the equation
-$\varphi^{\prime}(z)=\frac{\pi(z)}{\sigma(z)}\,$ (24)
+$\varphi^{\prime}(z)=\frac{\pi(z)}{\sigma(z)}\,,$ (26)
 where $\pi(z)$ is a polynomial at most of the first degree. In this case for $y(z)$ we obtain also a generalised hypergeometric-type equation:
@@
 $y_{n}(z)=\frac{B_{n}}{\rho(z)}\left[\sigma^{n}(z)\rho(z)\right]^{(n)},$ (37)
-where $B_{n}$ is some (normalization) constant, and the weight functions $\rho(z)$ satisfies Pearson’s equation Ismail
+where $B_{n}$ is some (normalization) constant, and the weight function $\rho(z)$ satisfies Pearson’s equation Ismail
 $(\sigma\rho)^{\prime}=\rho\tau\,.$ (38)

### v4

@@
 The derivative-of-the-Gaussian profile proposed by Gibbons and Hawking is approximated by the hyperbolic Scarf potential, which allows for an exact analytic solution via the Nikiforov-Uvarov method.
-Our results confirm the prediction of Zel’dovich and Polnarev about certain “magical” amplitudes of the potential.
+Our results confirm the prediction of Zel’dovich and Polnarev when the wave amplitude takes certain “magical” values.
 Key words: gravitational waves; displacement memory effect; flyby; Nikiforov-Uvarov method.
@@
 04.20.-q Classical general relativity;
-preprint: arXiv: 2502.0136v3 [gr-qc]
 I Introduction
@@
 Initial studies Ehlers considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories.
-Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : …although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.
+Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : “…although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.”
 The GW memory effect offers a way to test general relativity including the nonlinear nature of gravity, and provides additional information about the GW source.
@@
 This lead them to propose the derivative of the Gaussian,
-${\cal A}^{dG}\propto\frac{\;d}{dU}\Big{(}\exp[-U^{2}]\Big{)}\,.$ (2)
-as profile.
-The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
+${\cal A}^{dG}\propto\frac{\;d}{dU}\Big{(}\exp[-U^{2}]\Big{)}\,$ (2)
+as profile. The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
 ${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,.$ (3)
@@
 In this note we present a third approximation of the flyby profile (2), namely by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was first introduced in Scarf to mimic the potential of a real crystal.
-It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt1 .,
+It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt86 ; Levai ; Scrf1 .
+,
 ${\cal A}^{Scarf}(U)=2g\frac{\;\;d}{dU}\left(\frac{1}{\cosh U}\right)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (4)
@@
 Our main result then says that for the profile (4) analytic solutions can be found which confirm the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above.
-Our clue is that Eqn.(5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its (non-normalizable) zero-energy ground state solutions DM-1 .
+Our clue is that the geodesic equation (5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its real (non-normalizable) zero-energy ground states DM-1 .
 II Analytic solution for flyby
-We first recall how things go for the simple Pöschl-Teller case. DM-1 ; DM-2 . Putting $t=\tanh U$, into eqn. (5) with profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ yields
-$(1-t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}-2t\frac{dX^{\pm}}{dt}\mp\frac{g}{4}X^{\pm}=0\,,$ (7)
-Let us consider first one of the components, say $X^{+}$. Then DM is obtained in terms of Legendre polynomials DM-1 when the coefficient in the linear-in-$X^{+}$ term is
+We first recall how things go for the simple Pöschl-Teller case DM-1 ; DM-2 . Putting $t=\tanh U$ into eqn. (5) with profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ yields,
+$(1-t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}-2t\frac{dX^{\pm}}{dt}\mp\frac{g}{4}X^{\pm}=0\,.$ (7)
+Let us consider first one of the components, say $X^{+}$. Then DM is obtained in terms of Legendre polynomials DM-1 when the coefficient in the linear-in-$X^{+}$ term is quantized, as
 $g=-4n(n+1){\quad\text{\small where}\quad}n>0\quad\text{is a natural integer}\,.$ (8)
-However passing to the other component, the linear-in-$X^{-}$ term has opposite sign, therefore DM would be obtained for $-g$.
+However passing to the other component, the last, linear-in-$X^{-}$ term has opposite sign, therefore DM would be obtained for $-g$.
 But this is impossible because $g$ has already been fixed.
 In conclusion, one has what we called “half DM” DM-2 : we get a DM solution in one or in the other sector, but not for both in general — unless the “wrong” component is set to zero.
-Solutions for the Scarf potential ${\cal A}^{Scarf}$ (and similarly for derived Pöschl-Teller , ${\cal A}^{dPT}$), DM-2 ), are obtained by putting $t=\sinh{U}$. Then (5) is written
+Solutions for the Scarf potential ${\cal A}^{Scarf}$ (and similarly for derived Pöschl-Teller , ${\cal A}^{dPT}$, DM-2 ), are obtained by putting $t=\sinh{U}$. Then (5) is written
 $(1+t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}+t\frac{dX^{\pm}}{dt}{\mp}g\frac{t}{1+t^{2}}X^{\pm}=0\,.$ (9)
-The two equations here are again uncoupled and can therefore be solved independently.
-Note however the coefficient $t$ in front of the linear-in-$X^{\pm}$ term which compensates the $\pm$ sign change between the components.
+The two equations here are uncoupled and can therefore be solved independently.
+Note however the coefficient $t$ in front of the linear-in-$X^{\pm}$ term whose sign change compensates the $\pm$ sign change between the components.
 Therefore the same $g$ works for both components, as it will be seen in sec.
 III.
-Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller .
-For derived Pöschl-Teller it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions in eqn.
+Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller  ; it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions, see eqn.
 # (IV.16) of DM-2 .
 Turning to Scarf, we note that (9) is of the generalised hypergeometric type which can be solved analytically by the Nikiforov-Uvarov (NU) method NikiUvar .
-Since a fairly detailed description of this simple yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , here we just list the main result.
+Since a fairly detailed description of this simple yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , here we just list the main results.
 The algorithm, whose details are discussed in sec.III, yields the desired “magic” coefficient $g$ in (9),
 $|g|=|g_{n}|=(2n+1)\sqrt{n(n+1)}\,,$ (10)
-where $n$ is an integer. Then the trajectory is,
+where $n$ is a natural integer. Then the trajectory is,
 $\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,,}\;\;$ (11)
 where $y_{n}(t)$ is the polynomial
-$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,}\;\;$ (12)
+$y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,$ (12)
 with weight function
@@
 For $n\geq 5$, though, the numerical procedure becomes tedious as it requires to calculate a high number of decimals.
-For high wave numbers the result becomes also inaccurate because the procedure used in DM-2 is only an approximation.
-Plotting instead the analytic solution (11)-(12), FIG.3, is straightforward, underlining the advantage of the analytical solution.
+For high wave numbers the result becomes inaccurate because the procedure used in DM-2 is only an approximation.
+Plotting instead the analytic solution (11)-(12) yields accurate figures, underlining the advantage of the analytical approach.
 Figure 2: The DM geodesics for the derived Pöschl-Teller (3) and for the Scarf profiles (4) are similar however not fully identical, as shown for wave numbers ${\bf n=1}$ and ${\bf n=2}$.
 (The trajectories were rescaled for better comparison).
-Figure 3: DM geodesics for the Scarf profile (4) with wave numbers ${\bf n=5}$ and ${\bf n=6}$.
 III Transverse geodesics by the Nikiforov-Uvarov method
@@
 $\tau=\pi_{1}+2\pi=2t\pm\frac{|g|}{\sqrt{k}}\left|t+\frac{2k}{g}\right|.$ (18)
-To have real solutions, we assume $k>0$, that is we choose th upper sign $k=\frac{\sqrt{1+16g^{2}}-1}{8}$ in Eq.(16). On the other side, the quantization condition Eq. (36) requires
+To have real solutions, we assume $k>0$, that is we choose the upper sign $k=\frac{\sqrt{1+16g^{2}}-1}{8}$ in Eq.(16). On the other side, the quantization condition Eq. (37) requires
 $\lambda=-n(n+1)\mp\epsilon\,\frac{n|g|}{\sqrt{k}}\,,$ (19)
@@
 (10), for which we have polynomial solutions $y_{n}(t)$.
-The $y_{n}(t)$ can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
-The first three of them are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$
-It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. We have,
+There remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. We have,
 $\pi(t)=-nt-\mathrm{sign}(g)\sqrt{n(n+1)},\;\;\;\;\tau(t)=-(2n-1)t-2\mathrm{sign}(g)\sqrt{n(n+1)}\,.$ (20)
@@
 $\frac{d\varphi}{dt}=-\frac{nt+\mathrm{sign}(g)\sqrt{n(n+1)}}{1+t^{2}}\,,$ (21)
-can be easily integrated to yield (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
+can be easily integrated to yield (up to an irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
 $\varphi(t)=-\frac{n}{2}\ln(1+t^{2})+\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}\,.$ (22)
@@
 (11) implies the parity behaviour
-$X^{\pm}(-t)=(-1)^{n}X^{\mp}(t)\,$ (23)
-mentioned in sec.II. This curious pairing of the components is reminiscent of and is indeed closely related to supersymmetry MSUSY .
+$X^{\pm}_{n}(-t)=(-1)^{n}X^{\mp}_{n}(t)\,$ (23)
+mentioned in sec.II.
+Numerical evidence suggests that $t=0$ is not a root of $X^{\pm}_{n}$ neither for dPT nor for Scarf and, in fact, for a profile which is an order-odd derivative of an even function.
+On the contrary, it is a root when the profile is an order-even derivative.
+The curious pairing of the components in (23) is reminiscent of and is indeed closely related to supersymmetry MSUSY .
+The $y_{n}(t)$ can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
+The first three of them are $y_{1}(t)\sim-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$
 IV Conclusion
@@
 They are constructed by the Nikiforov-Uvarov method.
-The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.3, extending the observation made in DM-2 for the Gaussian and Pöschl-Teller profiles.
-The origin of half-waves is clear: their number is that of the zeros of the polynomial $y_{n}(t)$.
+The displacement effect (DM) arises when the wave zone contains $n$ half-waves as observed made DM-2 for the Gaussian and Pöschl-Teller profiles.
+The number of half-waves is equal to that of the zeros of the polynomial $y_{n}(t)$.
 Our Scarf solution is similar to but slightly different from those found semi-analytically for derived-Pöschl-Teller DM-2 , as seen in FIG.2.
@@
 Raposo .
-They include supersymmetry Scarfpot ; Scrf1 ; Compean ; Gendenshtein ; MSUSY and non-central potentials Dutt .
-The recent paper MSUSY , which is directly relevant for our purposes here, highlights the prominent role of SUSY for gravitational waves with Scarf profile.
-See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for further related work.
+They include supersymmetry Scarfpot ; Scrf1 ; Compean ; Gendenshtein ; MSUSY and non-central potentials Dutt86 .
+The recent paper MSUSY , which is directly relevant for our purposes here, highlights the prominent role of SUSY QM for gravitational waves with Scarf profile.
+See also Benin ; Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for further related work.
 We mention that our main tool here, namely the Nikiforov-Uvarov method, applies also to a number of other singular and periodic GW profiles, such as
@@
 Interestingly, the last profile corresponds to the Coulomb potential in constant curvature space $\mathbb{S}^{3}$, and hydrogenlike atoms in constant curvature spaces are most naturally considered by the Nikiforov-Uvarov method Alizzi .
+It has long been known (see, e.g., Gendenshtein ; Dutt86 ; Scarfpot ; Cooper94 ; Dutt92 ) that the Pöschl-Teller and (hyperbolic) Scarf potentials are parts of the same supersymmetric partner potentials, following from the shape invariant superpotential
+$W(U)=A{\,}\tanh(U)+B{\,}{\rm sech}(U)\,.$ (25)
+A recent paper MSUSY confirms our results using supersymmetric quantum mechanics with this superpotential.
 Acknowledgements.

### v5

@@
 Abstract
-We describe analytically the scattering of particles by a sandwich gravitational wave generated during a flyby.
-The derivative-of-the-Gaussian profile proposed by Gibbons and Hawking is approximated by the hyperbolic Scarf potential, which allows for an exact analytic solution via the Nikiforov-Uvarov method.
-Our results confirm the prediction of Zel’dovich and Polnarev when the wave amplitude takes certain “magical” values.
+The motion of particles hit by a burst of gravitational waves generated by flyby admits, for the derivative-of-the-Gaussian profile, only a numerical description.
+The profile can however be approximated by the hyperbolic Scarf potential which admits an exact analytic solution via the Nikiforov-Uvarov method.
+Our toy model is consistent with the prediction of Zel’dovich and Polnarev provided the wave amplitude takes certain “magical” values.
+Phys. Lett. B 868 (2025) 139687.  https://doi.org/10.1016/j.physletb.2025.139687
 Key words: gravitational waves; displacement memory effect; flyby; Nikiforov-Uvarov method.
@@
 04.20.-q Classical general relativity;
+preprint: arXiv: 2502.01326v5 [gr-qc]
 I Introduction
@@
 They argued indeed that Burst With Memory could be among the earliest kinds of gravitational waves detected BraTho .
-Initial studies Ehlers considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories.
+Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 .
+Initial studies Ehlers ; GWGNature considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories.
 Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : “…although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.”
@@
 Preliminary studies have suggested that the detection of GW memory from individual mergers of massive binary black holes is expected with LISA Favata .
+See Blanchet for a state-of-the-art review.
 However, more realistic estimates show that observing the GW memory effect at detectors such as LISA, although feasible, is a challenging task fly-1 ; fly-2 .
-Recent attempts to detect the effect in LIGO/Virgo’s first transient gravitational-wave catalog, as well as in a pulsar timing array experiment NANOGrav, have predictably yielded negative results fly-3 ; fly-4 .
+Recent attempts to detect the effect in LIGO/Virgo’s first transient gravitational-wave catalog, as well as in a pulsar timing array experiment NANOGrav, have, predictably, yielded negative results fly-3 ; fly-4 .
 However the prospects for the next generation of gravitational wave detectors are brighter even for future ground-based detectors such as Cosmic Explorer and the Einstein Telescope (see fly-2 ; fly-4 and references therein).
-Zel’dovich and Polnarev did not present any proof.
-Although their suggestion could be confirmed in Christo ; BMS1 ; BMS2 ; BMS3 ; BMS4 , there were some doubts due to the apparent contradiction with the Bondi-Pirani theorem on the caustic property of plane gravitational waves BoPi89 ; ShortMemory .
-Then it came as a surprise that that the DM effect could be confirmed numerically DM-1 ; Jibril ; DM-2 for plane waves — provided the amplitude takes some “magical” values.
-Below we present an analytic proof.
+Although the displacement effect could be confirmed Christo ; BMS1 ; BMS2 ; BMS3 ; BMS4 , there remained some doubts due to the apparent contradiction with the Bondi-Pirani theorem on the caustic property of plane gravitational waves BoPi89 ; ShortMemory .
+More recently, the DM effect could be verified numerically for plane waves using the approach proposed in LongMemory , provided the amplitude takes some “magical” values DM-1 ; Jibril ; DM-2 .
+In this note we study the Zel’dovich - Polnarev statement by using the simplified framework in ref LongMemory .
+Although they should be viewed only as toy models, they do provide some insight into the memory effect.
+Based Gibbons and Hawking’s earlier work Gibbons_1976 , the model in LongMemory proposes that the Brinkmann profile of the gravitational wave generated by flyby be approximated by the first derivative of the Gaussian (dG), eqn.
+(2) below.
+This yields however Sturm-Liouville-type geodesic equations with no analytic solution.
+In this paper we propose a similar but slightly different approximation, namely by the Scarf potential Scarf ; Scarfpot which does provide us with analytic solutions.
+The results are in full agreement with ZelPol .
 In Brinkmann coordinates, the metric is written as Brinkmann ,
-$ds^{2}=d{\bm{X}}^{2}+2dUdV+{\cal A}(U)\big{(}X_{+}^{2}-X_{-}^{2}\big{)}dU^{2}\,.$ (1)
-Gibbons and Hawking GibbHaw71 argued that in linear theory the Riemann tensor of a gravitational wave is proportional to the fourth time derivative of the quadrupole moment of the source GibbonsP .
-This lead them to propose the derivative of the Gaussian,
+$ds^{2}=d{\bm{X}}^{2}+2dUdV+{\cal A}(U)\big{(}X_{+}^{2}-X_{-}^{2}\big{)}dU^{2}\,,$ (1)
+where ${\bm{X}}$ is referred to as transverse coordinate; $U$ and $V$ are lightlike.
+In the non-relativistic Kaluza-Klein-type framework of ref.
+DBKP ; DGH91 $U$ plays the role of non-relativistic time.
+Ref.
+LongMemory argues that in linear theory the Riemann tensor of a gravitational wave is proportional to the fourth time derivative of the quadrupole moment of the source Gibbons_1976 ; GibbonsP .
+This lead us to propose the derivative of the Gaussian,
 ${\cal A}^{dG}\propto\frac{\;d}{dU}\Big{(}\exp[-U^{2}]\Big{)}\,$ (2)
-as profile. The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
+as profile. The geodesic equations for (2) have no analytic solution, though, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
 ${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,.$ (3)
@@
 Our investigations led us to conjecture that DM arises when a quantization holds, namely when the wave zone contains an integer number of what we called half-waves.
-In this note we present a third approximation of the flyby profile (2), namely by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was first introduced in Scarf to mimic the potential of a real crystal.
-It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt86 ; Levai ; Scrf1 .
+In this note we approximate the flyby profile (2) by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was initially introduced in Scarf to mimic the potential of a real crystal.
+It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt86 ; Levai ; Scrf1 ; MSUSY .
 ,
@@
 to be compared with (3).
+The difference is merely the exponent in the denominator.
 For small values of $U$ the Scarf profile (4) behaves as the derivative of the Gaussian or of Pöschl-Teller, (2) or (3), but for large $U$ it falls off more slowly.
@@
 We note also all three profiles in (2)-(3)-(4) are derivatives of an even function and are therefore odd.
-Figure 1: The derived Pöschl-Teller  profile (3) and its Scarf potential alter-ego (4) are both good approximations of the derived-Gaussian one proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
+Figure 1: The derived Pöschl-Teller  profile (3) and its Scarf potential alter-ego (4) are both approximations of the derived-Gaussian one proposed in LongMemory to describe flyby.
 The geodesics are, in Brinkmann coordinates, determined by the transverse equations DM-2 ,
-$\dfrac{d^{2}\!X^{\pm}}{dU^{2}}\mp\frac{1}{2}{\cal A}X^{\pm}=0\,,$ (5)
+$\dfrac{d^{2}\!X^{\pm}}{dU^{2}}\mp\frac{1}{2}{\cal A}(U)X^{\pm}=0\,,$ (5)
 where the $\pm$ label refers to the upper resp.
@@
 The sign of the second  term comes from the relative minus between the coordinates in (1).
-The equation for the $V$ component follows from these DM-1 ; DM-2 , and is therefore omitted.
-We recall that $U$ behaves, in the Bargmann framework DBKP ; DGH91 , as non-relativistic time.
 These Sturm-Liouville-type equations are completed with the DM boundary conditions,
 $\frac{d{\bm{X}}}{dU}\Big{|}(U=-\infty)=0=\frac{d{\bm{X}}}{dU}\Big{|}(U=+\infty)\,.$ (6)
-Our main result then says that for the profile (4) analytic solutions can be found which confirm the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above.
+Our main result then says that for the profile (4) analytic solutions can be found, consistently with the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above.
 Our clue is that the geodesic equation (5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its real (non-normalizable) zero-energy ground states DM-1 .
@@
 III.
-Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller  ; it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions, see eqn.
+Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller  ; it can be obtained in terms of confluent Heun functions DM-2 with numerically calculated approximate expressions, see eqn.
 # (IV.16) of DM-2 .
@@
 where $n$ is a natural integer. Then the trajectory is,
-$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,,}\;\;$ (11)
+$X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,,$ (11)
 where $y_{n}(t)$ is the polynomial
@@
 Thus we have
-$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.\quad$ (17)
+$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2},\;\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.\>$ (17)
 Then
@@
 The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq.
-(38) which requires to calculate the weight function $\rho(t)$ in (13).
+(38), which requires to calculate the weight function $\rho(t)$ in (13).
 Then we end up with eqn.
@@
 IV Conclusion
-The profile (2) proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by both the derived-Pöschl-Teller (3) and by the Scarf potentials (4), as illustrated in FIG.1.
-The advantage of Scarf is that it supports analytic solutions, (11)-(12), labeled by an integer $n$, eqn (10).
-They are constructed by the Nikiforov-Uvarov method.
-The displacement effect (DM) arises when the wave zone contains $n$ half-waves as observed made DM-2 for the Gaussian and Pöschl-Teller profiles.
+The toy profile (2) proposed in LongMemory , which may or may not be similar to that of flyby encountered in nature, yields, for certain “magic” values of the wave parameter, a step-like, pure displacement behavior, consistently with the suggestions of Zel’dovich and Polnarev.
+The profile (2) can, moreover, be approximated by both the derived-Pöschl-Teller (3) and by the Scarf potentials (4).
+The latter supports analytic solutions, (11)-(12), constructed by the Nikiforov-Uvarov method and illustrated in FIG.1.
+The solutions are labeled by an integer $n$, cf.
+eqn (10).
+The displacement effect (DM) arises when the wave zone contains $n$ half-waves as observed in DM-2 for the Gaussian and Pöschl-Teller profiles.
 The number of half-waves is equal to that of the zeros of the polynomial $y_{n}(t)$.
-Our Scarf solution is similar to but slightly different from those found semi-analytically for derived-Pöschl-Teller DM-2 , as seen in FIG.2.
-It has also the advantage to be exact, while the one in DM-2 is only approximate, with error increasing for higher wave numbers.
+Our Scarf solution is similar to but slightly different from those found semi-analytically for derived-Pöschl-Teller DM-2 FIG.2.
+It is exact, while the one in DM-2 is only approximate, with error increasing for higher wave numbers.
 The behavior of both the Scarf (4) and of the derived-Pöschl-Teller (3) is explained by their parity behaviour, (23), which explains in turn why do we get DM for both components.
@@
 $W(U)=A{\,}\tanh(U)+B{\,}{\rm sech}(U)\,.$ (25)
-A recent paper MSUSY confirms our results using supersymmetric quantum mechanics with this superpotential.
+A recent paper MSUSY confirms our results using supersymmetric quantum mechanics.
 Acknowledgements.